OpenStudy (anonymous):

A game is played using one die. If the die is rolled and shows a 5, the player wins $7. If the die 10) shows any number other than 5, the player wins nothing. If there is a charge of $1 to play the game, what is the game's expected value? A) $0.17 B) -$1 C) -$0.17 D) $6.00

OpenStudy (anonymous):

Can you tell me the formula ?

OpenStudy (anonymous):

\[E(X)=\sum_{i=1}^{6}x_i\cdot p(x_i)=$0\cdot \frac{ 5 }{ 6 }+$7\cdot \frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

oops, the number atop the sum should be 2. there are only 2 values, $0 and $7.

OpenStudy (anonymous):

Ok thank you